from vest, a jeepney accelerates uniformly over a time of 3.25 seconds and covers a distance of 15m. determine the acceleration of the jeepney.
Answer by Alan3354(69443)

from vest, a jeepney accelerates uniformly over a time of 3.25 seconds and covers a distance of 15m. determine the acceleration of the jeepney.
--------------
s = at^2/2
a = 2s/t^2 = 2*15/3.25^2
a =~ 2.84 m/sec/sec